# local non-bossiness

### v4

@@ 3. Local non-bossiness of DA
Given a school choice context $[N,S,\succ,q]$, let ${\rm{DA}}:{\mathcal{P}}\rightarrow{\mathcal{M}}$ be the student-optimal stable mechanism, namely the rule that associates to each school choice problem the outcome of the deferred acceptance algorithm when students make proposals.
@@
Given $[N,S,(\succ_{s},q_{s})_{s\in S},(P_{i})_{i\in N}]$,
Step 1: Each student $i$ proposes to the most preferred admissible school according to $P_{i}$, if any. Among those that have proposed to it, each school $s$ provisionally accepts the $q_{s}$-highest ranked students according to $\succ_{s}$.
Step $k\geq 2$: Each student $i$ who has not been provisionally accepted in step $k-1$ proposes to the most preferred admissible school according to $P_{i}$, among those to which she has not previously proposed, if any. Each school $s$ provisionally accepts the $q_{s}$-highest ranked students according to $\succ_{s}$, among those who have proposed to it in this step or were provisionally accepted by it in step $k-1$.
@@
For any preference profile $P\in{\mathcal{P}}$, the matching ${\rm{DA}}(P)$ assigns to each student the most preferred alternative in $S\cup\{s_{0}\}$ that she can reach in a stable matching under $P$ (Gale and Shapley, 1962; Roth, 1985). Moreover, the mechanism ${\rm{DA}}$ is strategy-proof (Dubins and Freedman, 1981; Roth, 1982).
However, when DA is implemented a student can modify the assignment of others without affecting her own (Ergin, 2002). Hence, one may wonder if there is a limit on the impact of these changes. In particular, under ${\rm{DA}}$, is it possible for a student to affect the set of students assigned to her school by reporting different preferences, without changing her school ? As the following theorem shows, the answer is negative.
@@
To get an intuitive idea of the proof of Theorem 1, let $\mu={\rm{DA}}(P)$ and $\mu^{\prime}={\rm{DA}}(P^{\prime}_{i},P_{-i})$ be distinct matchings that give the same assignment to student $i$, namely $\overline{s}\equiv\mu(i)=\mu^{\prime}(i)$.
By appealing to the Rural Hospital Theorem, we can show that students who receive different assignments in $\mu$ and $\mu^{\prime}$ are in schools that fill their quotas in both matchings. Hence, when we compare $\mu^{\prime}$ to $\mu$, each student who changes assignments displaces another. These movements induce cycles of students $(i_{1},\ldots,i_{r})$, where each $i_{k}$ receives in $\mu^{\prime}$ the school that $i_{k+1}$ had in $\mu$ (modulo $r$). If there exists a student who prefers $\mu^{\prime}$ to $\mu$, [^10] If this does not occur, we can maintain our arguments by simply swapping the roles of $P$ and $(P^{\prime}_{i},P_{-i})$. the stability of $\mu$ and $\mu^{\prime}$ ensures that at least one of these cycles is a $\mu$-improving cycle—that is, a cycle $(i_{1},\ldots,i_{r})$ in which every $i_{k}$ prefers $\mu(i_{k+1})$ to $\mu(i_{k})$ (modulo $r$).
We define a graph $G$ by letting each student $j$ who prefers $\mu^{\prime}$ to $\mu$ point to those who are assigned in $\mu$ to the school $\mu^{\prime}(j)$. Previous arguments ensure that $G$ has a $\mu$-improving cycle. Since $\mu$ is the student-optimal stable matching under $P$, any $\mu$-improving cycle necessarily induces justified envy. By allowing students in $G$ to block the improving cycles contained in it, we construct a new $\mu$-improving cycle $(i^{*}_{1},\ldots,i^{*}_{r})$ in which justified envy is limited to students outside $G$—that is, to those who are assigned in $\mu$ to schools that, in $\mu^{\prime}$, do not receive anyone who prefers $\mu^{\prime}$ to $\mu$ [^11] The $\mu$-improving cycle $(i^{*}_{1},\ldots,i^{*}_{r})$ is not necessarily present in $G$. Moreover, since students in $N\setminus\{i\}$ do not change their preferences, we can show that only $i$ blocks $(i^{*}_{1},\ldots,i^{*}_{r})$. Hence, the students in $\mu^{\prime}(\overline{s})$ weakly prefer $\mu$ to $\mu^{\prime}$. This result implies that either $\mu^{\prime}(\overline{s})\subseteq\mu(\overline{s})$ or $\mu(\overline{s})\subseteq\mu^{\prime}(\overline{s})$. Finally, to conclude that $\mu(\overline{s})=\mu^{\prime}(\overline{s})$, we invoke the Rural Hospital Theorem to ensure that any school with different entering classes in $\mu$ and $\mu^{\prime}$ always fills its quota.
@@
Fix a school choice context $[S,N,\succ,q]$ and w.l.o.g consider student $1$ and preference profiles $(P_{1},P_{-1}),(P^{\prime}_{1},P_{-1})\in{\mathcal{P}}$ such that $\overline{s}\equiv{\rm{DA}}_{1}(P_{1},P_{-1})={\rm{DA}}_{1}(P^{\prime}_{1},P_{-1})$ and ${\rm{DA}}(P_{1},P_{-1})\neq{\rm{DA}}(P^{\prime}_{1},P_{-1})$. We will prove that ${\rm{DA}}_{\overline{s}}(P_{1},P_{-1})={\rm{DA}}_{\overline{s}}(P^{\prime}_{1},P_{-1})$.
When $\overline{s}=s_{0}$, the result follows from the fact that no student who is not assigned a seat in ${\rm{DA}}(P_{1},P_{-1})$ can be assigned a seat in ${\rm{DA}}(P^{\prime}_{1},P_{-1})$ (see Lemma 4 in Appendix B).
Assume that $\overline{s}\neq s_{0}$. Let $\mu={\rm{DA}}(P_{1},P_{-1})$ and $\mu^{\prime}={\rm{DA}}(P^{\prime}_{1},P_{-1})$. Notice that, without loss of generality, we can assume that the set $I$ of students who prefer $\mu^{\prime}$ to $\mu$ is non-empty. [^12] Since $\mu(1)=\mu^{\prime}(1)$ and $\mu\neq\mu^{\prime}$, when $I$ is an empty set there is at least one student who prefers $\mu$ to $\mu^{\prime}$. Hence, our arguments hold by swapping the roles of $(P_{1},P_{-1})$ and $(P^{\prime}_{1},P_{-1})$.
Let $G=(V,E)$ be the graph with nodes $V=I\,\cup\,\{j:\mu(j)=\mu^{\prime}(i)\,\,\mbox{for some}\,\,i\in I\}$ and directed edges $E=\{[i,j]:i\in I,\,\mu(j)=\mu^{\prime}(i)\}.$ Hence, the nodes of $G$ are the students who prefer $\mu^{\prime}$ to $\mu$, together with those who are assigned in $\mu$ to the schools to which students in $I$ are assigned in $\mu^{\prime}$. Also, there is an edge between every $i\in I$ and all the students who are assigned to $\mu^{\prime}(i)$ in $\mu$.
@@
    A $\mu$-improving cycle is a tuple of different students $(i_{1},\ldots,i_{r})$ such that
@@
    A student $k$ blocks a $\mu$-improving cycle $(i_{1},\ldots,i_{r})$ when she $\mu$-blocks $[i_{l},i_{l+1}]$ for some $l\in\{1,\ldots,r\}$ [modulo $r$].
The following properties are satisfied:
  • (i)
    Each $i\in I$ is part of a $\mu$-improving cycle contained in $G$ (see Lemma 5 in Appendix B).
  • (ii)
    Each node in $G$ has a positive in-degree.
    By construction, each node in $V\setminus I$ has a positive in-degree. Moreover, the property (i) ensures that each student $i\in I$ is part of a $\mu$-improving cycle contained in $G$. Hence, there exists $j\in V$ such that $[j,i]\in E$.
Let $G^{\prime}=(V,E^{\prime})$ be the multigraph obtained from $G$ through the following edge-replacement procedure: given a student $j\in N$, substitute each edge $[i,j]\in E$ for which
$I[i,j]\equiv\{k\in V:\,\mbox{$k$ $\mu$-blocks $[i,j]$}\}\neq\emptyset$
by the directed edge $[\overline{k},j]$, where $\overline{k}$ is the student with the highest priority at school $\mu(j)$ among those in $I[i,j]$. Notice that, when $[\overline{k},j]\notin E$, its inclusion in $E^{\prime}$ is feasible because $\overline{k}\in I[i,j]\subseteq V$.
We make the following remarks regarding graph $G^{\prime}$.
@@
    The graph $G^{\prime}$ has a cycle $(i^{*}_{1},\ldots,i^{*}_{r})$.
    It follows from (ii) that all vertices in $V$ have a positive in-degree in $G^{\prime}$. Indeed, when $E^{\prime}$ is constructed from $E$, if some directed edge $[i,j]$ is deleted, then an edge who points to $j$ is included. As a consequence, $G^{\prime}$ has a cycle $(i^{*}_{1},\ldots,i^{*}_{r})$.
@@
    Any cycle in $G^{\prime}$ is a $\mu$-improving cycle.
    Remember that, if a student $k$ $\mu$-blocks $[i,j]\in E$, then $\mu(j)P_{k}\mu(k)$. Hence, each $[k,j]\in E^{\prime}\setminus E$ satisfies $\mu(j)P_{k}\mu(k)$. Since the definition of $E$ ensure that any $[k,j]\in E^{\prime}\cap E$ satisfies $\mu(j)P_{k}\mu(k)$, we conclude that any cycle in $G^{\prime}$ is a $\mu$-improving cycle.
@@
    Any cycle in $G^{\prime}$ cannot be blocked by any student in $V$.
@@
    The $\mu$-improving cycle $(i^{*}_{1},\ldots,i^{*}_{r})$ must be blocked by student $1$.
    Since $\mu$ is the student-optimal stable matching under $(P_{1},P_{-1})$, the $\mu$-improving cycle $(i^{*}_{1},\ldots,i^{*}_{r})$ must be blocked by some student in $N\setminus V$.
    For each $[i,j]\in E^{\prime}$, we have that $[i,j]\in E$ or $[i,j]\in E^{\prime}\setminus E$. Hence, $[i,j]\in E$ or there exists $[z,j]\in E$ that is $\mu$-blocked by $i$. This implies that either $\mu(j)=\mu^{\prime}(i)$ or there exists $z\in I$ such that [$\mu(j)=\mu^{\prime}(z)$, $\mu(j)P_{i}\mu(i)$, and $i\succ_{\mu(j)}z$]. As a consequence, when a student $k$ $\mu$-blocks an edge $[i,j]\in E^{\prime}$, we have that either [$\mu^{\prime}(i)P_{k}\mu(k)$ and $k\succ_{\mu^{\prime}(i)}i$] or [$\mu^{\prime}(z)P_{k}\mu(k)$ and $k\succ_{\mu^{\prime}(z)}z$] .
    We claim that no student $k\neq 1$ can block an edge in $E^{\prime}$. By contradiction, suppose that $k\neq 1$ $\mu$-blocks an edge $[i,j]\in E^{\prime}$. It follows from the arguments above and the stability of $\mu^{\prime}$ that there are two possible scenarios:
      If $\mu^{\prime}(i)P_{k}\mu(k)$ and $k\succ_{\mu^{\prime}(i)}i$, then $\mu^{\prime}(k)R_{k}\mu^{\prime}(i)$. Since $\mu^{\prime}(i)P_{k}\mu(k)$, the transitivity of $P_{k}$ ensures that $k\in I$. Hence, $k\in V$, which contradicts the property (3).
      If $\mu^{\prime}(z)P_{k}\mu(k)$ and $k\succ_{\mu^{\prime}(z)}z$, then $\mu^{\prime}(k)R_{k}\mu^{\prime}(z)$. Since $\mu^{\prime}(z)P_{k}\mu(k)$, the transitivity of $P_{k}$ ensures that $k\in I$. Hence, $k\in V$, which contradicts the property (3).
    We conclude that $(i^{*}_{1},\ldots,i^{*}_{r})$ must be blocked by student $1$.
We finish the proof by showing that $\mu(\overline{s})=\mu^{\prime}(\overline{s})$.
Since $(i^{*}_{1},\ldots,i^{*}_{r})$ is blocked by student $1$, we have that $1\notin V$. Thus, the school $\overline{s}$ does not receive in $\mu^{\prime}$ students who prefer $\mu^{\prime}$ to $\mu$.
We claim that $\mu^{\prime}(\overline{s})\subseteq\mu(\overline{s})$ or $\mu(\overline{s})\subseteq\mu^{\prime}(\overline{s})$. Indeed, there are two possible scenarios:
    If no one prefers $\mu$ to $\mu^{\prime}$, then everyone in $\mu^{\prime}(\overline{s})$ is indifferent between $\mu^{\prime}$ to $\mu$, which implies that $\mu^{\prime}(\overline{s})\subseteq\mu(\overline{s})$.
    If at least one student prefers $\mu$ to $\mu^{\prime}$, then we can repeat all the previous arguments swapping the roles of $(P_{1},P_{-1})$ and $(P^{\prime}_{1},P_{-1})$ in order to show that $\overline{s}$ does not receive in $\mu$ students who prefers $\mu$ to $\mu^{\prime}$. Notice that, if there exists $i\in\mu(\overline{s})\setminus\mu^{\prime}(\overline{s})$, then $i$ belongs to $I$. Hence, the property (i) guarantees that $i$ is part of a $\mu$-improving cycle contained in $G$, which ensures that there is $k\in I$ such that $[k,i]\in E$. This implies that $1\in V$, because $\mu^{\prime}(k)=\mu(i)=\overline{s}=\mu(1)$. A contradiction. Hence, $\mu(\overline{s})\setminus\mu^{\prime}(\overline{s})$ is an empty set, which implies that $\mu(\overline{s})\subseteq\mu^{\prime}(\overline{s})$.
Finally, since Lemma 4 guarantees that $|\mu(s)|=|\mu^{\prime}(s)|$ for all $s\in S$ such that $\mu(s)\neq\mu^{\prime}(s)$, we conclude that $\mu(\overline{s})=\mu^{\prime}(\overline{s})$. ∎
@@
As Theorem 1 demonstrates, when the ${\rm{DA}}$ mechanism is implemented, no student can modify her preferences to change her schoolmates without changing her school. But what happens when multiple students, all assigned to the same school, change their preferences? Can they change their colleagues without changing their school? The answer is negative, and not only for ${\rm{DA}}$, but also for any locally non-bossy and strategy-proof mechanism (see Lemma 2).
@@
Remark 1. Let $[N,S,\succ,q]$ be such that $|N|=\sum_{s\in S}q_{s}$, and assume that every student considers all schools admissible. We claim that when DA is implemented, a student who misreports her preferences and changes her assignment, also modifies all her colleagues. More formally, if ${\mathcal{P}}^{*}\subseteq{\mathcal{P}}$ is the preference domain in which everyone considers all schools admissible and ${\rm{Coll}}_{i}(P)=\{j\in N:j\neq i,\,\,{\rm{DA}}_{j}(P)={\rm{DA}}_{i}(P)\}$ is the set of colleagues of student $i$ in ${\rm{DA}}(P)$, then the following property holds (see Lemma 6):
  • For all $i\in N$, $P\in{\mathcal{P}}^{*}$, and $P^{\prime}_{i}\in{\mathcal{L}}$ such that $(P^{\prime}_{i},P_{-i})\in{\mathcal{P}}^{*}$,

### v5

@@ 3. Local non-bossiness of DA
Given a school choice context $N,S,\succ,q]$, let ${\rm{DA}}:{\mathcal{P}}\rightarrow{\mathcal{M}}$ be the student-optimal stable mechanism, namely the rule that associates to each school choice problem the outcome of the deferred acceptance algorithm when students make proposals.
@@
Given $N,S,(\succ_{s},q_{s})_{s\in S},(P_{i})_{i\in N}]$,
Step 1: Each student $i$ proposes to the most preferred admissible school according to $P_{i}$, if any. Among those who have proposed to it, each school $s$ provisionally accepts the $q_{s}$-highest ranked students according to $\succ_{s}$.
Step $k\geq 2$: Each student $i$ who has not been provisionally accepted in step k1k-1 proposes to the most preferred admissible school according to $P_{i}$, among those to which she has not previously proposed, if any. Each school $s$ provisionally accepts the $q_{s}$-highest ranked students according to $\succ_{s}$, among those who have proposed to it in this step or were provisionally accepted by it in step k1k-1.
@@
For every preference profile $P\in{\mathcal{P}}$, the matching ${\rm{DA}}(P)$ assigns to each student the most preferred alternative in $S\cup\{s_{0}\}$ that she can reach in a stable matching under $P$ (Gale and Shapley, 1962; Roth, 1985). Moreover, the mechanism $\rm{DA}$} is strategy-proof (Dubins and Freedman, 1981; Roth, 1982).
However, when DA is implemented, a student can modify the assignment of others without affecting her own (Ergin, 2002). Hence, one may wonder if there is a limit to the impact of these changes. In particular, under $\rm{DA}$}, is it possible for a student to affect the set of students assigned to her school by reporting different preferences, without changing her school ? As the following theorem shows, the answer is negative.
@@
To get an intuitive idea of the proof of Theorem 1, let $\mu={\rm{DA}}(P)$ and $\mu^{\prime}={\rm{DA}}(P^{\prime}_{1},P_{-1})$ be different matchings that give the same assignment to student $1$, namely $\overline{s}\equiv\mu(1)=\mu^{\prime}(1)$.
First, consider the benchmark case in which $P^{\prime}_{1}$ is a monotonic transformation of $P_{1}$ for $\overline{s}$, in the sense that the set of schools preferred to $\overline{s}$ shrinks when moving from $P_{1}$ to $P^{\prime}_{1}$. This assumption implies that $\mu$ is stable under $(P^{\prime}_{1},P_{-1})$, because if there were a blocking pair of $\mu$ under $(P^{\prime}_{1},P_{-1})$, it would be a blocking pair under $P$. Hence, as $\mu^{\prime}$ is the student-optimal stable matching under $(P^{\prime}_{1},P_{-1})$, $\mu^{\prime}$ Pareto dominates $\mu$ under PP. [^10] This argument proves the positivity of $\rm{DA}$} (cf., Raghavan, 2020, Proposition 2): if a student $i$ reports a monotonic transformation of $P_{i}$ for ${\rm{DA}}_{i}(P_{i},P_{-i})$ such that ${\rm{DA}}_{i}(P^{\prime}_{i},P_{-i})={\rm{DA}}_{i}(P_{i},P_{-i})$, then those who modify their assignments are better off. This property and the stability of $\mu$ guarantee that each student $i$ who changes her assignment from $\mu$ to $\mu^{\prime}$ displaces someone who was assigned to $\mu^{\prime}(i)$ at $\mu$. These movements induce $\mu$-improving cycles $(i_{1},\ldots,i_{r})$, where each student $i_{k}$ strictly prefers $\mu(i_{k+1})$ to $\mu(i_{k})$ under $P_{k}$ [modulo $r$]. Evidently, the implementation of a $\mu$-improving cycle induces justified envy under $P$, because $\mu$ is the student-optimal stable matching. We use this last property to ensure that $\overline{s}$ has the same assignment in $\mu$ and $\mu^{\prime}$. More precisely, denote by $V$ the set of students who are matched in $\mu$ to schools with different assignments in $\mu$ and $\mu^{\prime}$. Notice that $V$ includes the students who prefer $\mu^{\prime}$ to $\mu$. Let $G$ be the graph with nodes $V$ in which every $i$ that prefers $\mu^{\prime}$ to $\mu$ points to all students who are assigned in $\mu$ to $\mu^{\prime}(i)$. By allowing students in $V$ to block the changes from $\mu$ to $\mu^{\prime}$ that induce justified envy, in the proof of Theorem 1 we perform an edge replacement in $G$ in order to construct a $\mu$-improving cycle $(i^{*}_{1},\ldots,i^{*}_{r})$ in which justified envy is confined to students outside VV. [^11] Specifically, if the change of some student $i$ from $\mu(i)$ to $\mu^{\prime}(i)$ induces justified envy, we replace every edge from $i$ to a student $j$ such that $\mu(j)=\mu^{\prime}(i)$ by an edge from $k$ to $j$, where $k$ is the student with justified envy that has the highest priority at $\mu^{\prime}(i)$. In the proof of Theorem 1, we show that the multigraph obtained by these procedure has a $\mu$-improving cycle that no student in $V$ blocks. Since $\rm{DA}$} is stable and students in $N\setminus\{1\}$ have the same preferences in $P$ and $(P^{\prime}_{1},P_{-1})$, no student in $(N\setminus V)\setminus\{1\}$ blocks $(i^{*}_{1},\ldots,i^{*}_{r})$. Therefore, student $1$ must block $(i^{*}_{1},\ldots,i^{*}_{r})$, which implies that she does not belong to $V$. We conclude that $\mu(\overline{s})=\mu^{\prime}(\overline{s})$.
For an arbitrary $P_{1}^{\prime}$, let $P^{*}_{1}$ be a preference such that $\bar{s}$ is the best alternative. Given that $1$ is assigned to $\bar{s}$ in $\mu$ and $\mu^{\prime}$, we have that $\mu(1)={\rm{DA}}_{1}(P^{*}_{1},P_{-1})$ and $\mu^{\prime}(1)={\rm{DA}}_{1}(P^{*}_{1},P_{-1})$. Then, by the first argument, $\mu(\overline{s})={\rm{DA}}_{\overline{s}}(P^{*}_{1},P_{-1})$ and $\mu^{\prime}(\overline{s})={\rm{DA}}_{\overline{s}}(P^{*}_{1},P_{-1})$, which implies that $\mu(\overline{s})=\mu^{\prime}(\overline{s})$.
@@
Fix a school choice context $S,N,\succ,q]$ and without loss of generality consider student 11 and preference profiles $(P_{1},P_{-1}),(P^{\prime}_{1},P_{-1})\in{\mathcal{P}}$ such that $\mu\equiv{\rm{DA}}(P_{1},P_{-1})$ and $\mu^{\prime}\equiv{\rm{DA}}(P^{\prime}_{1},P_{-1})$ differ, but $\overline{s}\equiv\mu(1)$ is equal to $\mu^{\prime}(1)$. We will prove that $\mu(\overline{s})=\mu^{\prime}(\overline{s})$.
Case I. The preference relation $P^{\prime}_{1}$ is a monotonic transformation of $P_{1}$ for $\overline{s}$.
In this scenario, as $\{s\in S:sP^{\prime}_{1}\overline{s}\}\subseteq\{s\in S:sP_{1}\overline{s}\}$, $\mu$ is stable under $(P^{\prime}_{1},P_{-1})$. Since $\mu^{\prime}$ is the student-optimal stable matching under $(P^{\prime}_{1},P_{-1})$, $\mu^{\prime}(i)P_{i}\mu(i)$ for each $i\in N$ such that $\mu(i)\neq\mu^{\prime}(i)$. Let $I$ be the non-empty set of students who prefer $\mu^{\prime}$ to $\mu$. The stability of $\mu$ under $(P_{1},P_{-1})$ ensures that each student in $I$ is assigned in $\mu^{\prime}$ to a school that fills its quota at $\mu$. Hence, when we move from $\mu$ to $\mu^{\prime}$, each student in $I$ displaces another student.
Let $G=(V,E)$ be the graph with nodes and directed edges
$V=\bigcup_{\begin{subarray}{c}s\in S\cup\{s_{0}\}:\ \mu(s)\neq\mu^{\prime}(s)\end{subarray}}\mu(s),\quad\quad\quad\quad E=\{[i,j]:i\in I,\,\mu(j)=\mu^{\prime}(i)\}.$
Notice that $V$ is non-empty, as $I\subseteq V$. The nodes of $G$ are the students who are assigned in $\mu$ to schools that have different assignments in $\mu$ and $\mu^{\prime}$. Also, there is an edge between every $i\in I$ and all the students who are assigned to $\mu^{\prime}(i)$ in $\mu$.
@@
    A $\mu$-improving cycle [^12] We will also refer to a $\mu$-improving cycle as an improvement cycle from $\mu$. is a tuple of different students $(i_{1},\ldots,i_{r})$ such that
@@
    A student $k$ blocks a $\mu$-improving cycle $(i_{1},\ldots,i_{r})$ when she $\mu$-blocks $i_{l},i_{l+1}]$ for some $l\in\{1,\ldots,r\}$ [modulo $r$].
Let $G^{\prime}=(V,E^{\prime})$ be the multigraph obtained from $G$ through the following edge-replacement procedure: substitute each edge $i,j]\in E$ for which
$I[i,j]\equiv\{k\in V:\,k\ \ \mbox{$\mu$-blocks}\,\,[i,j]\}\neq\emptyset$
by the directed edge $\overline{k},j]$, where $\overline{k}$ is the student with the highest priority at school $\mu(j)$ among those in $I[i,j]$. Notice that, $I[i,j]=I[i,j^{\prime}]$ for all $i,j],[i,j^{\prime}]\in E$. Moreover, when $\overline{k},j]\notin E$, its inclusion in $E^{\prime}$ is feasible because $\overline{k}\in I[i,j]\subseteq V$.
We make the following remarks regarding $G^{\prime}$:
@@
    The graph $G^{\prime}$ has a cycle. That is, there exists a tuple of different students $(i^{*}_{1},\ldots,i^{*}_{r})$ such that $i^{*}_{l},i^{*}_{l+1}]\in E^{\prime}$ for all $l\in\{1,\ldots,r\}\,\,{\mbox{[modulo $r$]}}$.
    Remember that, when we move from $\mu$ to $\mu^{\prime}$, every student $i$ such that $\mu(i)\neq\mu^{\prime}(i)$ displaces someone who was assigned to $\mu^{\prime}(i)$ at $\mu$. Thus, by construction, all nodes of $G$ have a positive in-degree. The same property holds in $G^{\prime}$. Indeed, when $E^{\prime}$ is constructed from $E$, if some directed edge $[i,j]$ is deleted, then an edge that points to $j$ is included. Therefore, starting from a node in $V$ and moving backward through the edges in $E^{\prime}$ we can find a cycle of $G^{\prime}$.
@@
    Every cycle in $G^{\prime}$ is a $\mu$-improving cycle.
    Remember that, if a student $k$ $\mu$-blocks $i,j]\in E$, then $\mu(j)P_{k}\mu(k)$. Hence, each $k,j]\in E^{\prime}\setminus E$ satisfies $\mu(j)P_{k}\mu(k)$. Since the definition of $E$ ensures that every $k,j]\in E^{\prime}\cap E$ satisfies $\mu(j)P_{k}\mu(k)$, we conclude that each cycle in $G^{\prime}$ is a $\mu$-improving cycle.
@@
    A $\mu$-improving cycle in $G^{\prime}$ cannot be blocked by a student in $V$.
@@
    Every $\mu$-improving cycle $(i^{*}_{1},\ldots,i^{*}_{r})$ in $G^{\prime}$ is blocked by student $1$.
    Since $\mu$ is the student-optimal stable matching under $(P_{1},P_{-1})$, it follows from the previous remark that $(i^{*}_{1},\ldots,i^{*}_{r})$ must be blocked by some student in $N\setminus V$. Moreover, the stability of $\mu^{\prime}$ under $(P^{\prime}_{1},P_{-1})$ guarantees that only students in $I\cup\{1\}$ can block $(i^{*}_{1},\ldots,i^{*}_{r})$. Indeed, students who are not in $I\cup\{1\}$ have the same preferences in $P$ and $(P^{\prime}_{1},P_{-1})$, and stay in the same school under $\mu$ and $\mu^{\prime}$. If such a student blocks the $\mu$-improving cycle, she will also block $\mu^{\prime}$ because she will have more priority than one of the incoming students at some school. [^13] Assume that $h\neq 1$ blocks $(i^{*}_{1},\ldots,i^{*}_{r})$. Then, $h$ $\mu$-blocks some $k,j]\equiv[i^{*}_{l},i^{*}_{l+1}]$, where $l\in\{1,\ldots,r\}$ [modulo $r$]. If $s\equiv\mu(j)$, it follows that $sP_{h}\mu(h)$ and $h\succ_{s}k$. We have two possible scenarios: (i) if $k,j]\in E$, then $\mu^{\prime}(k)=s$; (ii) if $k,j]\in E^{\prime}\setminus E$, then there exists $i\in I$ such that $k$ $\mu$-blocks $i,j]\in E$, which ensures that $k\succ_{s}i$ and $\mu^{\prime}(i)=s$. Hence, the transitivity of $\succ_{s}$ guarantees that $h$ has more priority at $s$ than some student assigned to it at $\mu^{\prime}$. Therefore, as we assume that $sP_{h}\mu(h)$, if $\mu^{\prime}(h)=\mu(h)$, we would conclude that $(h,s)$ blocks $\mu^{\prime}$. As $I\subseteq V$, we conclude that $(i^{*}_{1},\ldots,i^{*}_{r})$ is blocked by student $1$.
It follows from remarks (3)-(4) that $1\notin V$, which implies that $\mu(\overline{s})=\mu^{\prime}(\overline{s})$. [^14] The construction of $G^{\prime}$ is not superfluous, because the $\mu$-improving cycle $(i^{*}_{1},\ldots,i^{*}_{r})$, that allows us to show that $\mu(\overline{s})=\mu^{\prime}(\overline{s}),$ may not be present in the graph $G$ (see 11).
Case II. The preference relation $P^{\prime}_{1}$ is arbitrary.
Let $P^{*}_{1}\in{\mathcal{L}}$ be such that $\bar{s}$ is the best alternative (it is possible that $P^{*}_{1}=P_{1}$). Given that $\mu$ is stable under $(P^{*}_{1},P_{-1})$, and ${\rm{DA}}(P^{*}_{1},P_{-1})$ is the student optimal stable matching under $(P^{*}_{1},P_{-1})$, we have that ${\rm{DA}}_{1}(P^{*}_{1},P_{-1})=\overline{s}$. Then, $\mu(1)=\mu^{\prime}(1)={\rm{DA}}_{1}(P^{*}_{1},P_{-1})=\overline{s}$. It follows from Case I that $\mu(\bar{s})={\rm{DA}}_{\overline{s}}(P^{*}_{1},P_{-1})$, and that $\mu^{\prime}(\bar{s})={\rm{DA}}_{\overline{s}}(P^{*}_{1},P_{-1})$. Therefore $\mu(\bar{s})=\mu^{\prime}(\bar{s})$. ∎
The proof of Theorem 1 uses the concept of an improvement cycle: an ordered list of students who exchange schools in a way that makes all of them better off. This concept was previously studied by Kesten (2010) to analyze efficiency adjustments to DA. However, the improvement cycle constructed in the proof of Theorem 1—the one that is blocked only by the student who misreports her preferences—does not necessarily coincide with one of the improvement cycles implemented by the algorithm proposed by Kesten (2010) for mitigating inefficiencies (see 11 and the subsequent discussion). This is not an issue, since our goal is not to address DA’s inefficiencies.
@@
As Theorem 1 demonstrates, when $\rm{DA}$} is implemented, no student can modify her preferences to change her schoolmates without changing her school. But what happens when multiple students, all assigned to the same school, change their preferences? Can they change their colleagues without changing their school? The answer is negative, and not only for $\rm{DA}$}, but also for every locally non-bossy and strategy-proof mechanism (see Lemma 2).
Remark 1.
@@
Given $N,S,\succ,q]$, when DA is implemented, a student who misreports her preferences and changes her assignment, also modifies all of her colleagues. More formally, if ${\rm{Coll}}_{i}(P)=\{j\in N:j\neq i,\,\,{\rm{DA}}_{j}(P)={\rm{DA}}_{i}(P)\}$ is the set of colleagues of a student $i$ in ${\rm{DA}}(P)$, then the following property holds (see Lemma 7):
  • For all $i\in N$, $P\in{\mathcal{P}}$, and $P^{\prime}_{i}\in{\mathcal{L}}$,